# Farsight: Fostering Responsible AI Awareness During AI Application Prototyping

### v1

@@ 4.3. Harm Envisioner
Interactive Node-link Tree Visualization. After clicking the “Envision Consequences & Harms” button in the Awareness Sidebar, Harm Envisioner appears as a pop-up window on top of the prompt-crafting tool (Fig. 7). It begins with a text box filled with an LLM-generated summary of a user’s prompt (Fig. 7B). The user is prompted to revise the summary to align with the target task in their prompt. Next, the window transitions into an interactive node-link tree visualization (Reingold and Tilford, 1981), where the user can pan and zoom to navigate the view (Fig. 7C). First, the window shows the user’s prompt summary as the root of the tree which is visualized as a text box. The user can click the root node and the LLM will generate potential use cases of an AI application based on the user’s prompt, and the use cases are visualized as the root’s children nodes. Similarly, users can click a generated node and the LLM will generate its children nodes (stakeholders and then harms). There is a max of four layers, following an order of the user’s prompt summary $\rightarrow$ use cases $\rightarrow$ stakeholders $\rightarrow$ harms. This layer order reflects the recommended harm envisioning workflow in responsible AI literature (Doteveryone, 2019; Microsoft, 2022b; Barocas et al., 2021; Madaio et al., 2022; Microsoft, 2022a) and helps users to comprehend and organize diverse harms across different contexts (G2). For additional examples of LLM-generated use cases, stakeholders, and harms in Farsight, see Table S1.
@@
Fig. 8. Icons used to represent different harm types.
@@ 6.1. Participants
Software Engineer
3, 4, 5, 6, 7, 12, 13, 15, 16, 17, 19, 23, 25, 26, 28, 29, 33, 34, 35, 41, 42
Product Manager
1, 8, 10, 11, 14, 20, 24, 27, 36
Linguist
2, 21, 30, 31
AI Researcher
9, 18, 39, 40
UX Researcher
22
Data Scientist
32
Test Engineer
37
Marketing Specialist
38
@@ 6.2.4. Harm Rating
Our collected harms were either (1) directly envisioned by participants or (2) exported from Farsight or Farsight Lite and subsequently curated by participants during Harm Envisioning Activities and Harm Envisioning Activities. Each harm included the impacted stakeholders and a description of the harm. After removing duplicates and random shuffling, we randomly and evenly assigned harms to raters via spreadsheet format. Raters had access to the details of the intended AI feature of each harm, including the prompt and the context of the AI feature (e.g., the prompt and context in § C.1). To prevent the raters from being influenced by our hypotheses, we did not include the experimental conditions in the rating sheet. In other words, raters did not know if a harm was from a Farsight user, a Farsight Lite user, or a Envisioning Guide user. To mitigate rating noise, we designated three raters for each harm. As identifying likely and severe harms is often an objective in AI harm envisioning exercises (Microsoft, 2022a; Raji et al., 2020), we asked raters to rate each harm’s likelihood and severity on a 4-point Likert scale (strongly agree, agree, disagree, and strongly disagree to statements “This harm is likely to occur for this stakeholder” and “This harm will severely impact this stakeholder”). Raters could also choose an N/A option if they perceived a rating was not applicable for that feature or use case. During data analysis, we numericalized these four categories as ordinal scores: 1, 2, 3, 4 and removed N/As. See Table S2 for a random subset of harms that were collected from participants and their corresponding ratings.
@@ 6.4. Findings: Changes in Users’ Envisioning Ability and Approach (RQ1)
Failure-mode-driven envisioning
Participants envisioned harm by initially considering the AI feature’s failure modes (e.g., wrong summarization), limitations of LLMs (e.g., hallucination), or vulnerabilities within system implementation (e.g., data storage). This pattern is similar to a Failure Mode and Effects Analysis (Rismani et al., 2023).
Usage-driven envisioning
Participants envisioned harm by initially considering who may be impacted through this feature and in what usage scenario, such as students using the article summarizer for completing assignments. Then, participants envisioned potential harms that might impact the stakeholders within the identified scenario.
Consider high-stakes uses
Participants deliberately thought about high-stakes use cases of the AI feature, such as being used in medical, financial, and legal domains.
Consider misuses
Participants deliberately envisioned potential misuse of the AI feature, where malicious actors like scammers and hackers could exploit this AI feature to cause harm.
Consider indirect stakeholders
Participants deliberately brainstormed stakeholders indirectly impacted by the AI feature, such as people who did not use the AI tools, individuals mentioned in the input text, and society at large.
Consider cascading harms
Participants deliberately considered (1) harms that could result from other harms, such as productivity loss due to AI errors can lead to economic loss; or (2) harms that might occur even when the AI feature operated as expected, such as students using AI to cheat in homework.
@@ 7.2. Subjectivity in Harm Envisioning
In our evaluation user study, many participants report challenges overcoming the limitations of their own experiences and perspectives when envisioning harms (C2). In addition, we also observed the seven RAI raters of participants’ harms disagreed about which harms were more or less severe or likely, resulting in a low inter-rater reliability for these two dimensions (‡ E, Table S2). Our empirical findings contribute to prior research that highlights the role of subjectivity and positionality in anticipating harms (Boyarskaya et al., 2020; Liu et al., 2022) and in data annotation, particularly for annotations of toxicity or hate speech (e.g., Mostafazadeh Davani et al., 2022; Díaz et al., 2022; Denton et al., 2021; Davani et al., 2023; Prabhakaran et al., 2023). What constitutes harm and the assessment of harm severity are often influenced by the individual’s background, lived experiences, or even the organizational culture they are working in (Whittaker, 2021; of QueerInAI et al., 2023). For example, for the article summarizer (Harm Envisioning Activities), one participant envisioned a harm scenario: “If the summary is wrong, journalists’ reputation might be harmed.” (Table S2). This harm scenario received likelihood ratings of 1, 4, and 3, and severity ratings of 1, 3, and 4 from three randomly assigned raters. It is possible that the rater who assigned the ratings of 3 and 4 possessed specific knowledge about the harms of journalists using LLMs to write article summaries, which led them to rate this harm scenario as more likely and more severe.
@@ B.2. Example Farsight Output
This table shows the outputs of Farsight for a toxicity classifier prompt. Three columns are ”Use Cases”, ”Stakeholders”, and ”Harms.”
@@
Online moderators use it to remove toxic content from social media.  [Uncaptioned image]
Online moderators
Online moderators may be exposed to toxic content while using the AI tool.
@@
Social media users
Social media users may be unfairly censored by the AI tool.
@@
The company
The company may be sued for defamation if the AI tool mislabels a post as toxic.
@@
Social media users who do not use this AI product
Social media users who do not use this AI product may be exposed to toxic content.
@@
Social media advertisers
Social media advertisers may lose revenue due to toxic content being removed from social media.
@@
Customer service agents use it to identify abusive customers.  [Uncaptioned image]
Customer service agents
Customer service agents may feel stressed or anxious about the accuracy of the AI tool.
@@
Abusive customers
Abusive customers may be denied service due to being labeled as toxic.
@@
The company
The company may be sued for defamation if the AI tool mislabels a customer as abusive.
@@
Customers who are not abusive
Customers who are not abusive may be misidentified as abusive and treated poorly by customer service agents.
@@
Customers who are misidentified as abusive
Customers who are misidentified as abusive may be denied service or treated poorly by customer service agents.
@@
Law enforcement
Law enforcement may make false arrests or detentions of innocent people due to AI-generated toxicity labels.
@@
Law enforcement uses it to identify potential terrorists.  [Uncaptioned image]
Potential terrorists
Potential terrorists may be unfairly targeted by law enforcement due to AI-generated misclassifications.
@@
Victims of terrorism
Victims of terrorism may be re-traumatized by AI-generated content that glorifies or incites violence against them.
@@
Civil rights groups
Civil rights groups may be unfairly targeted by law enforcement due to AI-generated false positives.
@@
People who are falsely identified as potential terrorists
People who are falsely identified as potential terrorists may be subject to increased surveillance, harassment, and discrimination.
@@
Military uses it to identify potential insurgents.  [Uncaptioned image]
Military
Military may use AI to target and harm innocent civilians.
@@
Insurgency
Insurgency may be able to evade detection by using AI-generated text that is not classified as toxic.
@@
Governments
Governments may use AI to target and surveil innocent people.
@@
Civilians who are wrongly identified
Civilians who are wrongly identified may be subject to violence or discrimination.
@@
Civilians who are not identified as insurgents
Civilians who are not identified as insurgents may be subject to violence or discrimination.
@@
Hate groups
Hate groups may be able to recruit more people due to the AI tool.
@@
Potential recruits
Potential recruits may be exposed to toxic content that could lead to radicalization.
@@
Hate groups use it to identify potential recruits.  [Uncaptioned image]
The company
The company may be held liable for hate speech generated by its AI product.
@@
People who are targeted by hate groups
People who are targeted by hate groups may be harassed or threatened by hate groups.
@@
People who are not targeted by hate groups
People who are not targeted by hate groups may be exposed to toxic content.
@@
Scammers use it to identify potential victims.  [Uncaptioned image]
Victims of scams
Victims of scams may lose money or personal information.
@@
Scammers
Scammers may be able to target more vulnerable people with their scams.
@@
The company
The company may be held liable for the harm caused by scammers using the AI tool.
@@
Victim’s financial institutions
Victim’s financial institutions may lose money due to scammers using AI to identify potential victims.
@@
Family and friends of the victim
Family and friends of the victim may be harmed by the scammer’s actions.
@@ Appendix E Harm Rating Processing and Inter-rater Reliability
In our evaluation user study, we recruited seven raters to rate the likelihood and severity of each collected harm (§ 6.2.4). In total, we collected 989 harms with 895 unique harms (Table S2). We randomly assign each unique harm to three raters to rate its likelihood and severity on a 4-point Likert scale (strongly agree, agree, disagree, and strongly disagree to statements “this harm is likely to happen” and “this harm is severe”). Raters could also choose an N/A option if they perceived a rating was not applicable. After collecting all ratings, we dropped all N/As and numericalized four rating categories as ordinal scores: 1, 2, 3, 4.
@@ E.1. Example Harms Collected from Participants
Email summarizer (Harm Envisioning Activities)
AI company
AI company’s reputation may be harmed due to inaccurate summary
3, 3, 3
2, 2, 3
Author
@@
The AI summary might lose context and information, critical context. It could cut out some crucial part of the email.
4, 3, 3
3, 3, 3
Owner of the email
@@
Rewrite could leak data from other users’ email.
3, 1, 2
4, 4, 2
People whose names / personal information is lost in the summarization
@@
Potential data loss. The email could be very long, and the rewrite could miss some context (e.g., full name). The rewrite can cause confusion to the receivers.
3, 3, 3
2, 3, 3
Product team, company
Prompt injection attack.
3, 1, 3
2, 1, 3
Receiver
@@
Omission of vital information and context. The feature can miss some very important information.
3, 3, 3
3, 2, 3
Receiver
@@
The initial rewrite is a bit open ended, which might cause anxiety.
2, 3, 2
1, 3, 2
Recipients
@@
If users can control the temperature of the LLM, they can generate more inaccurate emails.
1, 3, 1
4, 2, 1
@@
Recipients due to loss of productivity and insight into their orgs metrics
The email may miss some metrics from the email.
3, 3, 3
2, 2, 3
Scam victims
@@
Scam emails are very common. Mass emailing a lot of people with easy automation can cause harm to victims.
2, 3, 3
4, 3, 3
Sender
@@
In a professional setting, if the feature is not sufficient to capture the keywords, the summary can sound unprofessional. It can harm the reputation of the sender.
2, 3, 3
2, 2, 2
Sender
@@
The summary is inaccurate, e.g., having the wrong time, names in the rewrite. It can harm the sender’s reputation.
3, 3, 3
2, 3, 3
Sender, Recipient
@@
The feature introduces incorrect information that is not in the original email.
3, 3, 3
3, 2, 2
Sender, recipient
@@
Quality of the communication. If the tone of the rewrite is different from the original email, it can cause harm to both sender and recipient. Miscommunication.
3, 2, 3
2, 2, 3
Sender, recipient
@@
The delicacy in the original email may be removed in the shorter email. The tone, emotion, and human elements may be changed.
4, 3, 3
3, 2, 3
Sender, recipient
@@
The summary can have a different voice / style from the email (more directly, less soft). It can be perceived as negative.
4, 3, 4
3, 2, 4
Sender, recipient, company relationships
@@
AI can be culturally blind. It might not follow some typical email norm in the rewrite.
3, 3, 3
2, 2, 3
Senders, receivers
@@
The AI feature puts words on other people’s behalf. It can have different viewpoints from the email. It can have misrepresented views.
2, 3, 4
3, 3, 4
Senders, recipient
@@
The summary can lose some tone, voice of the email, becoming less personal.
3, 3, 4
2, 3, 3
Stakeholders of the decision
@@
If the communication is about decision making and the rewrite contains error, it can harm the stakeholders of the decision.
1, 4, 3
1, 4, 3
Team building the feature
@@
Damage to reputation, credibility to have a feature fail this way.
3, 2, 3
3, 1, 2
Toxicity classifier (Harm Envisioning Activities)
Advertisers
@@
If the AI feature fails, ads might be present next to toxic content, losing revenues.
3, 3, 3
1, 2, 3
Bullies
@@
Teachers use it to identify students being bullied. The AI may make the teachers be biased against the bullies.
2, 3, 2
3, 3, 2
Company of the AI product
The company may lose users due to the presence of more moderation.
3, 2, 1
4, 2, 1
Customer
@@
Customer service agents use it to identify abusive customers. Customers may lose the opportunity to get help from customer service agents.
3, 3, 3
3, 2, 4
Customer service agents
@@
If customer service uses it to screen people, false positives: exposing customer service agents to toxic customers.
3, 1, 3
3, 1, 3
Employees of the company
@@
HR departments use it to screen job applicants for toxic behavior. Employees may be unfairly rejected for jobs due to AI-generated toxicity labels.
1, 3, 3
1, 3, 4
End user (children & parents)
@@
Mislabeled outputs will not be blocked by parental controls.
3, 3, 3
3, 4, 4
Government
@@
Military uses it to identify potential insurgents. Governments may use AI to target and surveil innocent people, leading to human rights violations.
3, 2, 3
4, 3, 3
Hate group targets
@@
Hate groups use it to identify potential recruits. Hate groups may be able to recruit more people due to the AI tool.
3, 2, 3
3, 3, 3
Job applicants
@@
Reddit has a karma score. Similarly, if a social media uses this feature to prioritize non-toxic content, misclassification can lose job opportunities for social media users (e..g, tweet not seen by companies).
3, 4, 1
3, 3, 1
Online moderators
@@
Online moderators may lose their jobs because the AI tool can already distinguish toxic vs non-toxic so there is no need for a moderator.
2, 1, 3
3, 1, 4
Online moderators
@@
Online moderators may miss some toxic content and post it on social media.
3, 3, 3
3, 4, 3
People being targeted by law enforcement
@@
Law enforcement uses it to identify potential terrorists. Potential terrorists ([different name]) may be unfairly targeted by law enforcement due to AI-generated misclassifications.
3, 1, 3
3, 1, 4
People who are not members of hate groups
@@
Hate groups use it to identify potential recruits. People who are not members of hate groups may be affected by increased hate group activity.
3, 3, 4
3, 3, 4
Sender, Receivers
@@
False positive. Flagging non-toxic content as toxic, causing harm to legitimate receivers of the information (filtered). Sender cannot post due to the filter.
3, 3, 4
3, 3, 4
Social media users
@@
Social media users may be exposed to toxic content that was mislabeled by the AI as non-toxic.
3, 3, 3
3, 4, 3
Social media users
3, 3, 3
2, 4, 2
Social media users
Social media users who do not use this AI might be exposed to some distorted message.
2, 1, 3
3, 1, 3
Social media users
Unfairly flagged as toxic (false positive trigger).
3, 3, 3
3, 2, 2
Student

### v2

@@ 4.3. Harm Envisioner
Interactive Node-link Tree Visualization. After clicking the “Envision Consequences & Harms” button in the Awareness Sidebar, Harm Envisioner appears as a pop-up window on top of the prompt-crafting tool (Fig. 7). It begins with a text box filled with an LLM-generated summary of a user’s prompt (Fig. 7B). The user is prompted to revise the summary to align with the target task in their prompt. Next, the window transitions into an interactive node-link tree visualization (Reingold and Tilford, 1981), where the user can pan and zoom to navigate the view (Fig. 7C). First, the window shows the user’s prompt summary as the root of the tree which is visualized as a text box. The user can click the root node and the LLM will generate potential use cases of an AI application based on the user’s prompt, and the use cases are visualized as the root’s children nodes. Similarly, users can click a generated node and the LLM will generate its children nodes (stakeholders and then harms). There is a max of four layers, following an order of the user’s prompt summary $\rightarrow$ use cases $\rightarrow$ stakeholders $\rightarrow$ harms. This layer order reflects the recommended harm envisioning workflow in responsible AI literature (Doteveryone, 2019; Microsoft, 2022b; Barocas et al., 2021; Madaio et al., 2022; Microsoft, 2022a) and helps users to comprehend and organize diverse harms across different contexts (G2). For additional examples of LLM-generated use cases, stakeholders, and harms in Farsight, see LABEL:table:farsight-output.
@@ 6.1. Participants
Software Engineer 3, 4, 5, 6, 7, 12, 13, 15, 16, 17, 19, 23, 25, 26, 28, 29, 33, 34, 35, 41, 42
Product Manager 1, 8, 10, 11, 14, 20, 24, 27, 36
Linguist 2, 21, 30, 31
AI Researcher 9, 18, 39, 40
UX Researcher 22
Data Scientist 32
Test Engineer 37
Marketing Specialist 38
[…]
@@ 6.2.4. Harm Rating
Our collected harms were either (1) directly envisioned by participants or (2) exported from Farsight or Farsight Lite and subsequently curated by participants during Harm Envisioning Activities and Harm Envisioning Activities. Each harm included the impacted stakeholders and a description of the harm. After removing duplicates and random shuffling, we randomly and evenly assigned harms to raters via spreadsheet format. Raters had access to the details of the intended AI feature of each harm, including the prompt and the context of the AI feature (e.g., the prompt and context in § C.1). To prevent the raters from being influenced by our hypotheses, we did not include the experimental conditions in the rating sheet. In other words, raters did not know if a harm was from a Farsight user, a Farsight Lite user, or a Envisioning Guide user. To mitigate rating noise, we designated three raters for each harm. As identifying likely and severe harms is often an objective in AI harm envisioning exercises (Microsoft, 2022a; Raji et al., 2020), we asked raters to rate each harm’s likelihood and severity on a 4-point Likert scale (strongly agree, agree, disagree, and strongly disagree to statements “This harm is likely to occur for this stakeholder” and “This harm will severely impact this stakeholder”). Raters could also choose an N/A option if they perceived a rating was not applicable for that feature or use case. During data analysis, we numericalized these four categories as ordinal scores: 1, 2, 3, 4 and removed N/As. See LABEL:table:participant-harm for a random subset of harms that were collected from participants and their corresponding ratings.
@@ 6.4. Findings: Changes in Users’ Envisioning Ability and Approach (RQ1)
Failure-mode-driven envisioning Participants envisioned harm by initially considering the AI feature’s failure modes (e.g., wrong summarization), limitations of LLMs (e.g., hallucination), or vulnerabilities within system implementation (e.g., data storage). This pattern is similar to a Failure Mode and Effects Analysis (Rismani et al., 2023).
Usage-driven envisioning Participants envisioned harm by initially considering who may be impacted through this feature and in what usage scenario, such as students using the article summarizer for completing assignments. Then, participants envisioned potential harms that might impact the stakeholders within the identified scenario.
Consider high-stakes uses Participants deliberately thought about high-stakes use cases of the AI feature, such as being used in medical, financial, and legal domains.
Consider misuses Participants deliberately envisioned potential misuse of the AI feature, where malicious actors like scammers and hackers could exploit this AI feature to cause harm.
Consider indirect stakeholders Participants deliberately brainstormed stakeholders indirectly impacted by the AI feature, such as people who did not use the AI tools, individuals mentioned in the input text, and society at large.
Consider cascading harms Participants deliberately considered (1) harms that could result from other harms, such as productivity loss due to AI errors can lead to economic loss; or (2) harms that might occur even when the AI feature operated as expected, such as students using AI to cheat in homework.
[…]
@@ 7.2. Subjectivity in Harm Envisioning
In our evaluation user study, many participants report challenges overcoming the limitations of their own experiences and perspectives when envisioning harms (C2). In addition, we also observed the seven RAI raters of participants’ harms disagreed about which harms were more or less severe or likely, resulting in a low inter-rater reliability for these two dimensions (‡ E, LABEL:table:participant-harm). Our empirical findings contribute to prior research that highlights the role of subjectivity and positionality in anticipating harms (Boyarskaya et al., 2020; Liu et al., 2022) and in data annotation, particularly for annotations of toxicity or hate speech (e.g., Mostafazadeh Davani et al., 2022; Díaz et al., 2022; Denton et al., 2021; Davani et al., 2023; Prabhakaran et al., 2023). What constitutes harm and the assessment of harm severity are often influenced by the individual’s background, lived experiences, or even the organizational culture they are working in (Whittaker, 2021; of QueerInAI et al., 2023). For example, for the article summarizer (Harm Envisioning Activities), one participant envisioned a harm scenario: “If the summary is wrong, journalists’ reputation might be harmed.” (LABEL:table:participant-harm). This harm scenario received likelihood ratings of 1, 4, and 3, and severity ratings of 1, 3, and 4 from three randomly assigned raters. It is possible that the rater who assigned the ratings of 3 and 4 possessed specific knowledge about the harms of journalists using LLMs to write article summaries, which led them to rate this harm scenario as more likely and more severe.
@@ B.2. Example Farsight Output
Online moderators use it to remove toxic content from social media.  [Uncaptioned image] Online moderators Online moderators may be exposed to toxic content while using the AI tool.
@@
Social media users Social media users may be unfairly censored by the AI tool.
@@
The company The company may be sued for defamation if the AI tool mislabels a post as toxic.
@@
Social media users who do not use this AI product Social media users who do not use this AI product may be exposed to toxic content.
@@
Social media advertisers Social media advertisers may lose revenue due to toxic content being removed from social media.
@@
Customer service agents use it to identify abusive customers.  [Uncaptioned image] Customer service agents Customer service agents may feel stressed or anxious about the accuracy of the AI tool.
@@
Abusive customers Abusive customers may be denied service due to being labeled as toxic.
@@
The company The company may be sued for defamation if the AI tool mislabels a customer as abusive.
@@
Customers who are not abusive Customers who are not abusive may be misidentified as abusive and treated poorly by customer service agents.
@@
Customers who are misidentified as abusive Customers who are misidentified as abusive may be denied service or treated poorly by customer service agents.
@@
Law enforcement Law enforcement may make false arrests or detentions of innocent people due to AI-generated toxicity labels.
@@
Law enforcement uses it to identify potential terrorists.  [Uncaptioned image] Potential terrorists Potential terrorists may be unfairly targeted by law enforcement due to AI-generated misclassifications.
@@
Victims of terrorism Victims of terrorism may be re-traumatized by AI-generated content that glorifies or incites violence against them.
@@
Civil rights groups Civil rights groups may be unfairly targeted by law enforcement due to AI-generated false positives.
@@
People who are falsely identified as potential terrorists People who are falsely identified as potential terrorists may be subject to increased surveillance, harassment, and discrimination.
@@
Military uses it to identify potential insurgents.  [Uncaptioned image] Military Military may use AI to target and harm innocent civilians.
@@
Insurgency Insurgency may be able to evade detection by using AI-generated text that is not classified as toxic.
@@
Governments Governments may use AI to target and surveil innocent people.
@@
Civilians who are wrongly identified Civilians who are wrongly identified may be subject to violence or discrimination.
@@
Civilians who are not identified as insurgents Civilians who are not identified as insurgents may be subject to violence or discrimination.
@@
Hate groups Hate groups may be able to recruit more people due to the AI tool.
@@
Potential recruits Potential recruits may be exposed to toxic content that could lead to radicalization.
@@
Hate groups use it to identify potential recruits.  [Uncaptioned image] The company The company may be held liable for hate speech generated by its AI product.
@@
People who are targeted by hate groups People who are targeted by hate groups may be harassed or threatened by hate groups.
@@
People who are not targeted by hate groups People who are not targeted by hate groups may be exposed to toxic content.
@@
Scammers use it to identify potential victims.  [Uncaptioned image] Victims of scams Victims of scams may lose money or personal information.
@@
Scammers Scammers may be able to target more vulnerable people with their scams.
@@
The company The company may be held liable for the harm caused by scammers using the AI tool.
@@
Victim’s financial institutions Victim’s financial institutions may lose money due to scammers using AI to identify potential victims.
@@
Family and friends of the victim Family and friends of the victim may be harmed by the scammer’s actions.
@@ Appendix E Harm Rating Processing and Inter-rater Reliability
In our evaluation user study, we recruited seven raters to rate the likelihood and severity of each collected harm (§ 6.2.4). In total, we collected 989 harms with 895 unique harms (LABEL:table:participant-harm). We randomly assign each unique harm to three raters to rate its likelihood and severity on a 4-point Likert scale (strongly agree, agree, disagree, and strongly disagree to statements “this harm is likely to happen” and “this harm is severe”). Raters could also choose an N/A option if they perceived a rating was not applicable. After collecting all ratings, we dropped all N/As and numericalized four rating categories as ordinal scores: 1, 2, 3, 4.
@@ E.1. Example Harms Collected from Participants
Email summarizer (Harm Envisioning Activities) AI company AI company’s reputation may be harmed due to inaccurate summary 3, 3, 3 2, 2, 3
@@
Author The AI summary might lose context and information, critical context. It could cut out some crucial part of the email. 4, 3, 3 3, 3, 3
@@
Owner of the email Rewrite could leak data from other users’ email. 3, 1, 2 4, 4, 2
@@
People whose names / personal information is lost in the summarization Potential data loss. The email could be very long, and the rewrite could miss some context (e.g., full name). The rewrite can cause confusion to the receivers. 3, 3, 3 2, 3, 3
Product team, company Prompt injection attack. 3, 1, 3 2, 1, 3
@@
Receiver Omission of vital information and context. The feature can miss some very important information. 3, 3, 3 3, 2, 3
@@
Receiver The initial rewrite is a bit open ended, which might cause anxiety. 2, 3, 2 1, 3, 2
@@
Recipients If users can control the temperature of the LLM, they can generate more inaccurate emails. 1, 3, 1 4, 2, 1
@@
Recipients due to loss of productivity and insight into their orgs metrics The email may miss some metrics from the email. 3, 3, 3 2, 2, 3
@@
Scam victims Scam emails are very common. Mass emailing a lot of people with easy automation can cause harm to victims. 2, 3, 3 4, 3, 3
@@
Sender In a professional setting, if the feature is not sufficient to capture the keywords, the summary can sound unprofessional. It can harm the reputation of the sender. 2, 3, 3 2, 2, 2
@@
Sender The summary is inaccurate, e.g., having the wrong time, names in the rewrite. It can harm the sender’s reputation. 3, 3, 3 2, 3, 3
@@
Sender, Recipient The feature introduces incorrect information that is not in the original email. 3, 3, 3 3, 2, 2
@@
Sender, recipient Quality of the communication. If the tone of the rewrite is different from the original email, it can cause harm to both sender and recipient. Miscommunication. 3, 2, 3 2, 2, 3
@@
Sender, recipient The delicacy in the original email may be removed in the shorter email. The tone, emotion, and human elements may be changed. 4, 3, 3 3, 2, 3
@@
Sender, recipient The summary can have a different voice / style from the email (more directly, less soft). It can be perceived as negative. 4, 3, 4 3, 2, 4
@@
Sender, recipient, company relationships AI can be culturally blind. It might not follow some typical email norm in the rewrite. 3, 3, 3 2, 2, 3
@@
Senders, receivers The AI feature puts words on other people’s behalf. It can have different viewpoints from the email. It can have misrepresented views. 2, 3, 4 3, 3, 4
@@
Senders, recipient The summary can lose some tone, voice of the email, becoming less personal. 3, 3, 4 2, 3, 3
@@
Stakeholders of the decision If the communication is about decision making and the rewrite contains error, it can harm the stakeholders of the decision. 1, 4, 3 1, 4, 3
@@
Team building the feature Damage to reputation, credibility to have a feature fail this way. 3, 2, 3 3, 1, 2
@@
Toxicity classifier (Harm Envisioning Activities) Advertisers If the AI feature fails, ads might be present next to toxic content, losing revenues. 3, 3, 3 1, 2, 3
@@
Bullies Teachers use it to identify students being bullied. The AI may make the teachers be biased against the bullies. 2, 3, 2 3, 3, 2
Company of the AI product The company may lose users due to the presence of more moderation. 3, 2, 1 4, 2, 1
@@
Customer Customer service agents use it to identify abusive customers. Customers may lose the opportunity to get help from customer service agents. 3, 3, 3 3, 2, 4
@@
Customer service agents If customer service uses it to screen people, false positives: exposing customer service agents to toxic customers. 3, 1, 3 3, 1, 3
@@
Employees of the company HR departments use it to screen job applicants for toxic behavior. Employees may be unfairly rejected for jobs due to AI-generated toxicity labels. 1, 3, 3 1, 3, 4
@@
End user (children & parents) Mislabeled outputs will not be blocked by parental controls. 3, 3, 3 3, 4, 4
@@
Government Military uses it to identify potential insurgents. Governments may use AI to target and surveil innocent people, leading to human rights violations. 3, 2, 3 4, 3, 3
@@
Hate group targets Hate groups use it to identify potential recruits. Hate groups may be able to recruit more people due to the AI tool. 3, 2, 3 3, 3, 3
@@
Job applicants Reddit has a karma score. Similarly, if a social media uses this feature to prioritize non-toxic content, misclassification can lose job opportunities for social media users (e..g, tweet not seen by companies). 3, 4, 1 3, 3, 1
@@
Online moderators Online moderators may lose their jobs because the AI tool can already distinguish toxic vs non-toxic so there is no need for a moderator. 2, 1, 3 3, 1, 4
@@
Online moderators Online moderators may miss some toxic content and post it on social media. 3, 3, 3 3, 4, 3
@@
People being targeted by law enforcement Law enforcement uses it to identify potential terrorists. Potential terrorists ([different name]) may be unfairly targeted by law enforcement due to AI-generated misclassifications. 3, 1, 3 3, 1, 4
@@
People who are not members of hate groups Hate groups use it to identify potential recruits. People who are not members of hate groups may be affected by increased hate group activity. 3, 3, 4 3, 3, 4
@@
Sender, Receivers False positive. Flagging non-toxic content as toxic, causing harm to legitimate receivers of the information (filtered). Sender cannot post due to the filter. 3, 3, 4 3, 3, 4
@@
Social media users Social media users may be exposed to toxic content that was mislabeled by the AI as non-toxic. 3, 3, 3 3, 4, 3